# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v1

@@ Abstract
In this work, we propose Wasserstein distributionally robust shallow convex neural networks (WaDiRo-SCNNs) to provide reliable nonlinear predictions when subject to adverse and corrupted datasets. Our approach is based on a new convex training program for $\operatorname{ReLU}$ shallow neural networks which allows us to cast the problem as an exact, tractable reformulation of its order-1 Wasserstein distributionally robust equivalent. Our training procedure is conservative by design, has low stochasticity, is solvable with open-source solvers, and is scalable to large industrial deployments. We provide out-of-sample performance guarantees and show that hard convex physical constraints can be enforced in the training program. WaDiRo-SCNN aims to make neural networks safer for critical applications, such as in the energy sector. Finally, we numerically demonstrate the performance of our model on a synthetic experiment and a real-world power system application, i.e., the prediction of non-residential buildings’ hourly energy consumption. The experimental results are convincing and showcase the strengths of the proposed model.
@@ 1 Introduction
Neural networks (NNs) have been traditionally overlooked in critical sectors, e.g., energy [1], healthcare [2], and finance [3]. Even though they tend to be successful nonlinear predictors, their lack of interpretability, limited out-of-the-box performance guarantees, complex training procedures, as well as high susceptibility to data corruption and other adversarial attacks, have labelled them as unreliable for such applications. Many solutions have been proposed in the past decade to guide and frame their training as much as possible, e.g., blackbox optimization (BBO) algorithms with provable convergence properties for hyperparameter tuning [4, 5], tight MLOps (contraction of Machine Learning, software development and operations) procedures for lifecycle management [6], post-training verification frameworks [7, 8, 9], introduction of physical knowledge into the models [10, 11], or a combination of them [12]. These methods were pursued to make up for the fact that (i) neural networks generally have a non-convex loss function with respect to their weights as well as their hyperparameters [13, 14]; (ii) their high complexity and opacity make them unpredictable and theoretical verification is mandatory for complete performance assessment; (iii) it is hard to guarantee their out-of-sample performance over time as their drift needs to be monitored; (iv) their training often requires a large amount of data and physical insights greatly reduce the data requirements [15]. Nevertheless, these methods contribute to an even higher training complexity and do not scale easily for large deployments.
In this work, we are interested in neural networks with guarantees emerging from the training procedure itself. Our motivation is to develop a scalable nonlinear predictor with inherent theoretical guarantees and low-stochasticity training for large-scale critical applications, e.g., virtual power plants [16], where reliable forecasts of distributed assets are crucial for the safe operation of the electrical power grid.
@@
We specifically target the prediction of the localized impact of non-residential buildings enrolled in demand response programs, e.g., energy consumption or peak-shaving potential. Because of their size and diverse activities, non-residential buildings have complex consumption patterns that require prediction tools able to model such complexity, e.g., neural networks. As each building needs its own model, scalability issues arise when utilizing complex ML training procedures. Finally, the sensitivity of the application, and the risk associated with poor predictions, promote models with theoretical guarantees and robustness to possible anomalies in the training data, e.g., numerical errors, noise in sensor readings, telemetry issues, meters outage, and unusual extreme events [17].
@@ 1.1 Related Work
We now review the bodies of literature closest to ours.
Convex training of neural networks
@@
Convex optimization for machine learning (ML) training is desirable as a convex training program guarantees convergence to the minimal loss. In recent years, [18] have shown that training a feedforward shallow neural network (SNN), i.e., a one-hidden layer NN, with $\operatorname{ReLU}$ activation functions is equivalent to a convex training procedure. They also extended their results to convolutional neural networks. More importantly, they fundamentally linked $\operatorname{ReLU}$-SNN training problem to a regularized convex program. Following, [19] have obtained an efficient convex training procedure of SNNs with GPU acceleration, and [20] have designed an adversarial training by controlling the worst-case output of the neural network in a neighbourhood around each training samples. In this work, we generalize the regularization procedure of feedforward SNNs in regression tasks through the lens of distributionally robust optimization with the Wasserstein metric.
Distributionally robust neural networks
@@
Distributionally robust optimization (DRO), in ML, aims at minimizing the expected training loss of a model over the most adverse data distribution within a distance of the training samples. It minimizes the worst-case out-of-sample risk associated with the training of the model [21]. Different authors have worked on distributionally robust neural networks. Reference [22] have trained deep neural networks using the first-order approximation of the Wasserstein DRO loss function for backpropagation. The authors of [23] have developed a DRO gradient estimator that integrates into neural networks’ optimization pipeline similarly to the widely used stochastic gradient descent (SGD) optimizer. In [24], the authors have proposed a method to train deep NNs in a DRO fashion by defining the ambiguity set with groups of the training data and increasing regularization. These methods integrate distributional robustness in the training procedure of deep NNs but do not directly tackle the Wasserstein DRO problem, or an equivalent reformulation. Our proposition, while only applying to shallow networks, works with the tractable reformulation of the Wasserstein DRO problem proposed by [25].
Physics-informed neural networks
@@
The training approach for physics-informed neural networks primarily involves a weighted multi-objective function designed to narrow the gap between labels and model predictions while penalizing physical constraints violation. This is achieved by simultaneously minimizing the residuals of a partial differential equation (PDE) and its boundary conditions [15]. The loss terms related to the PDE are derived through the automatic differentiation of the neural network’s outputs [26]. This complex loss function is then progressively optimized through backpropagation within the network. While authors of [15] note that boundary conditions can be strictly imposed during training by forcing a part of the network to satisfy these conditions, see [27], they are typically integrated as penalty terms within the loss function. Physics-informed neural networks are characterized by their extensive modelling capabilities, facilitated by automatic differentiation, that allows for the approximation of continuous PDEs across various neural network architectures, e.g., CNNs [28], LSTMs [29], BNNs [30], and GANs [31]. Although limited to convex constraints, in this work, we demonstrate that our training methodology can enforce hard constraints on the shallow network’s output during training. These constraints may include, but are not limited to, boundary constraints, ramping constraints, and constraints derived from discrete-time dynamical systems.
Application
@@
As stated earlier, prediction algorithms used to forecast non-residential buildings’ consumption patterns must respect some inherent constraints. They need to be computationally lightweight, they must be able to model nonlinearities, and the sensibility of the application requires trustworthiness in their predictions, e.g., robustness, interpretability, and guarantees. Because grid operators do not generally have access to information on building activity, occupancy, and control, data-driven approaches tend to be the default option [32]. Authors have used support vector regression (SVR) [33] and other hybrid methods leveraging both classification and linear regression to satisfy the aforementioned constraints [32]. These methods are linear by parts and can approximate nonlinear distributions adequately. Statistical methods such as ARIMA [34] and Gaussian processes have also proved to be good contenders as they offer empirical uncertainty bounds [35]. Many authors have proposed deep-learning methods, yet these propositions do not have much industrial acceptability as they fail to meet the industrial constraints, and thus are not considered. In this work, we show that our proposition is interesting for virtual power plant applications where the forecast of many distributed assets is used for critical decision-making as we satisfy all industrial constraints better than other comparable models.

### v2

@@ Abstract
In this work, we propose Wasserstein distributionally robust shallow convex neural networks (WaDiRo-SCNNs) to provide reliable nonlinear predictions when subject to adverse and corrupted datasets. Our approach is based on a new convex training program for $\operatorname{ReLU}$-based shallow neural networks which allows us to cast the problem as an exact, tractable reformulation of its order-1 Wasserstein distributionally robust counterpart. Our training procedure is conservative, has low stochasticity, is solvable with open-source solvers, and is scalable to large industrial deployments. We provide out-of-sample performance guarantees, show that hard convex physical constraints can be enforced in the training program, and propose a mixed-integer convex post-training verification program to evaluate model stability. WaDiRo-SCNN aims to make neural networks safer for critical applications, such as in the energy sector. Finally, we numerically demonstrate the performance of our model on a synthetic experiment, a real-world power system application, i.e., the prediction of non-residential buildings’ hourly energy consumption in the context of virtual power plants, and on benchmark datasets. The experimental results are convincing and showcase the strengths of the proposed model.
@@ 1 Introduction
Consider the following general feedforward neural network (NN) training problem for a network of $h\in\mathbb{N}$ hidden layers under a loss function $\mathcal{L}:\mathbb{R}^{N\times m_{h+1}}\times\mathbb{R}^{N\times m_{h+1}}\to\mathbb{R}$ which penalizes the difference between the $N\in\mathbb{N}$ training labels and NN predictions:
$\min_{\mathbf{W}_{1},\mathbf{W}_{2},\ldots,\mathbf{W}_{h},\mathbf{b}_{1},\mathbf{b}_{2},\ldots,\mathbf{b}_{h}}$ $\mathcal{L}\left(\hat{\mathbf{Y}},\mathbf{Y}\right)$ (NNT)
s.t. $\hat{\mathbf{y}}_{j}=\phi^{h}(\mathbf{W}_{h}^{\top}(\ldots\phi^{2}(\mathbf{W}_{2}^{\top}\phi^{1}(\mathbf{W}_{1}^{\top}\mathbf{x}_{j}+\mathbf{b}_{1})+\mathbf{b}_{2})\ldots)+\mathbf{b}_{h})\quad$ $\forall j\in\llbracket N\rrbracket,$
where $\phi_{i}:\mathbb{R}^{m_{i}}\to\mathbb{R}^{m_{i+1}}$ is the activation function of the hidden layer $i$, $m_{i}\in\mathbb{N}^{*}$ is the number of neurons in layer $i$, $\mathbf{W}_{i}\in\mathbb{R}^{m_{i-1}\times m_{i}}$ are the weights in layer $i$, and $\mathbf{b}_{i}\in\mathbb{R}^{m_{i}}$ are the bias weights of layer $i$ for all $i\in\{1,2,3,\ldots,h\}=\llbracket h\rrbracket$. Let $\mathbf{x}_{j}\in\mathbb{R}^{m_{0}},\mathbf{y}_{j}\in\mathbb{R}^{m_{h+1}},\ \text{and}\ \hat{\mathbf{y}}_{j}\in\mathbb{R}^{m_{h+1}}$ define training features, training labels, and training predictions, respectively, $\forall j\in\llbracket N\rrbracket$. Matrices $\mathbf{Y},\hat{\mathbf{Y}}\in\mathbb{R}^{N\times m_{h+1}}$ are the collection of all labels and predictions. In our notation, we assume that NN inputs and outputs are vectors. A common activation function $\phi_{i}$ for regression is the element-wise rectified linear unit ($\operatorname{ReLU}$) operator:
$\operatorname{ReLU}(\cdot)=\max\{\mathbf{0},\cdot\},$
as it is accurate when modelling nonlinear patterns and biomimics the required excitation threshold for brain neurons to activate. The acquisition of features is done through the input layer and the output of labels through the output layer such that a neural network of $h$ hidden layers has effectively $h+2$ layers. To solve (NNT), the loss is propagated iteratively, one layer at a time, from the output neurons to the input neurons, by computing gradient estimators of the NN architecture at each layer and updating weights with methods derived from non-convex optimization, e.g., stochastic gradient descent (Amari 1993). This procedure is referred to as backpropagation (Bishop and Bishop 2023).
In part because (NNT) is typically non-convex and only solvable through stochastic heuristics with many hyperparameters, NNs have been traditionally disregarded in critical sectors, e.g., energy (Chatzivasileiadis et al. 2022), healthcare (Rasheed et al. 2022), and finance (Giudici and Raffinetti 2023). Even though they tend to be successful nonlinear predictors, their lack of interpretability, limited out-of-the-box performance guarantees, significant data requirements, as well as high susceptibility to data corruption and other adversarial attacks, have labelled them as unreliable for such applications.
In the past decade, many solutions have been proposed to guide and certify NN training as much as possible to mitigate these inherent flaws before deployment. For example, blackbox optimization (BBO) algorithms with provable convergence properties can add some local convergence guarantees in the hyperparameter tuning phase (Audet et al. 2022, Kawaguchi and Sun 2021) because NNs’ loss landscape, with respect to their hyperparameters, can be fractal (Sohl-Dickstein 2024). However, when paired with highly stochastic training programs, their relevance is diminished.
Because their high complexity and opacity make them unpredictable, complete performance assessment is hard to obtain during training. As such, post-training verification frameworks have emerged as a way to certify formal guarantees on trained machine learning (ML) models (Huang et al. 2017). These procedures are conceptually similar to quality control tests in factories. By using a model’s trained weights as parameters and representing the model’s internal mechanisms in an optimization problem, these frameworks aim to find theoretical worst-case scenarios, e.g., worst-case constraint violation in a specific critical application (Venzke et al. 2020), worst-case instability (Huang et al. 2021), and more (Albarghouthi 2021). A general feedforward neural network verification problem takes the form:
$\max_{\mathbf{x},\hat{\mathbf{y}},\bm{\sigma}}$ $\mathcal{V}\left(\hat{\mathbf{y}},\mathbf{x},\bm{\sigma}\right)$ (PTV)
s.t. $\hat{\mathbf{y}}=\phi^{h}(\mathbf{W}_{h}^{\top}(\ldots\phi^{2}(\mathbf{W}_{2}^{\top}\phi^{1}(\mathbf{W}_{1}^{\top}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2})\ldots)+\mathbf{b}_{h})$
$f_{i}(\hat{\mathbf{y}},\mathbf{x},\bm{\sigma})\leq 0$ $\forall i\in\llbracket n\rrbracket$
$g_{i}(\hat{\mathbf{y}},\mathbf{x},\bm{\sigma})=0$ $\forall i\in\llbracket l\rrbracket$
$\mathbf{x}\in\mathcal{X},$
where $\mathbf{x}\in\mathcal{X}$ is any possible input from the feature space $\mathcal{X}\subseteq\mathbb{R}^{m_{0}}$, $\mathcal{V}:\mathbb{R}\times\mathbb{R}^{m_{h+1}}\times\mathbb{R}^{\operatorname{card}(\bm{\sigma})}\to\mathbb{R}$ is an objective function which measures a specific guarantee, $\bm{\sigma}$ represents all of the problem’s other variables used in the verification process, and $f_{1},f_{2},\dots,f_{n},g_{1},g_{2},\dots,g_{l}$ are the family of inequality and equality constraints needed for the certification. Note that a convex representation of the internal mechanisms with respect to the inputs is extremely desirable for such problems because the convergence guarantees to global optimality of convex optimization eliminate any doubt on the verification’s validity.
When a neural network is used in a dynamical setting, i.e., a system that evolves with time over a state-space and follows a set of physical rules (Wiggins 2003), the introduction of physical knowledge from the setting into its training procedure can increase its overall reliability and greatly reduce the data requirements to obtain good empirical performances (Karniadakis et al. 2021, Xu et al. 2023). This type of training is often referred to as physics-informed, physics-guided, or physics-constrained depending on how physical knowledge is embedded in the training process. We refer interested readers to Cuomo et al. (2022) for an overview of different techniques. We will cover the most common one in Section 1.2.
Typically, as it is hard to guarantee NN out-of-sample performance over time and their performance may degrade (drift), the aforementioned techniques are combined in complex pre-deployment and post-deployment loops (Stiasny et al. 2022) with tight MLOps (contraction of machine learning, software development, and operations) procedures for lifecycle management (Kreuzberger et al. 2023). Good MLOps practices help monitor and assert, theoretically and empirically, the quality of ML models from training to deployment but require great computational infrastructure to be well implemented. Such infrastructures are not democratized in every sector, especially energy and healthcare. We present an example of a safe MLOps pipeline inspired by Stiasny et al. (2022) in Figure 1. A safe MLOps pipeline can be divided into three parts: (i) definition of initial conditions, (ii) hyperparameter optimization, and (iii) model verification. Initial conditions englobe the choice of dataset, ML model, hyperparameter tuning algorithm, and the definition of the hyperparameter search space. Hyperparameter optimization covers the typical training and validation phases: hyperparameters are chosen for each run to pre-process the training dataset and train the ML model, the trained model is then evaluated on the validation dataset, and different metrics are logged. This part ends when a stopping criterion is met, e.g., the maximum number of runs is reached. Finally, the best models in the validation phase are verified through post-training certification methods and evaluated on the testing dataset. If the performance expectations are not met by any model, initial conditions must be redefined. Otherwise, the best ML model can be deployed. Most of the computational burden stems from hyperparameter optimization. As such, there is a growing interest in designing complex predictive models with low-stochasticity training procedures, e.g., few hyperparameters and convex with respect to the weights, as well as lower data pre-processing requirements, e.g., by being distributionally robust. These models simplify the hyperparameter optimization phase and have more chance than standard models to meet the requirements of the verification phase.
Dataset Model selection Hyperparameter tuning algorithm and search space Feature selection Informed sampling Outlier filtering Scaling Data preprocessing Training Validation Log run Hyperparameter selection Stop? no yes Testing Theoretical certifications (iii) Model verification Good enough? no yes Deploy model Redefine initial conditions model val. loss new training hyperparameters best model in validation new data hyperparameters (ii) Hyperparameter optimization (i) Initial conditions
Figure 1: Example of safe pre-deployment MLOps pipeline
In this work, our motivation is to design a nonlinear predictive ML model with a simple low-stochasticity training program that can be guided by physical constraints and that has inherent performance guarantees. We also aim to design this model with an architecture that can be easily included in post-training verification frameworks or data-driven control applications, if needs be. Shallow convex neural networks, as we will show, are perfect candidates for the task. In doing so, we hope to lower computational infrastructure needs for safe deployment in large-scale critical sectors where a lack of expertise and budget constraints do not permit state-of-the-art MLOps pipelines at all stages, and yet accurate nonlinear predictions are required. We specifically target virtual power plants (VPPs) (Pudjianto et al. 2007) applications where reliable forecasts of distributed assets are crucial for the safe operations of the electrical power grid.
@@
As an example of a standard VPP application, we utilize our model to predict the hourly energy consumption of non-residential buildings enrolled in demand response programs (Siano 2014). Because of their immense impact on distribution grids and their diverse activities, non-residential buildings have complex nonlinear consumption patterns that must be predicted accurately to enhance grid operation (Massana et al. 2015). As each building needs its model and thus possibly thousands of models must be deployed, scalability issues arise when utilizing complex stochastic ML training procedures in complete safe MLOps pipelines. Finally, the sensitivity of the application, and the risk associated with poor predictions, promote models with theoretical guarantees, mechanisms to respect physical constraints, and robustness to possible anomalies in the training data, e.g., numerical errors, noise in sensor readings, telemetry issues, meters outage, and unusual extreme events (Zhang et al. 2021).
1.1 Notation
We use the following notation throughout the paper:
  • Bold lowercase letters refer to vectors, bold uppercase letters refer to matrices, and normal lowercase or uppercase letters refer to scalars. Bold Greek letters may refer to matrices depending on context.
  • The sets of real, natural, and integer numbers are denoted by $\mathbb{R}$, $\mathbb{N}$, and $\mathbb{Z}$, respectively.
  • Calligraphic uppercase letters refer to sets, e.g., $\mathcal{X}$ and $\mathcal{Z}$, or functions depending on context, e.g., $\mathcal{L}(\cdot)$ and $\mathcal{V}(\cdot)$.
[…]
  • Let $\mathbf{a}\in\mathbb{R}^{n}$ be a general vector of dimension $n\in\mathbb{N}^{*}$, column vectors are compactly written $\mathbf{a}=\left(a_{1},a_{2},\ldots,a_{n}\right)$ and row vectors as $\mathbf{a}^{\top}=\left(a_{1},a_{2},\ldots,a_{n}\right)^{\top}=\left(a_{1}\ a_{2}\ \ldots\ a_{n}\right)$.
[…]
  • $\text{1}(\cdot):\mathbb{R}^{n}\to\{0,1\}^{n}$ is an element-wise indicator function of arbitrary dimension $n$ which returns 1 or 0 element-wise if the condition is true or false, respectively, for each separate element.
  • $\operatorname{vec}(\cdot):\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$ denotes the vectorization operator which flattens matrices. If $\mathbf{A}=(\mathbf{a}_{1}\ \mathbf{a_{2}}\ \ldots\ \mathbf{a_{n}})$ where $\mathbf{a}_{i}\in\mathbb{R}^{m},\ n,m\in\mathbb{N}^{*},\ \forall i\in\llbracket n\rrbracket$, then $\operatorname{vec}(\mathbf{A})=(\mathbf{a}_{1}^{\top}\ \mathbf{a}_{2}^{\top}\ \ldots\ \mathbf{a}_{n}^{\top})^{\top}$.
[…]
  • $\mathbf{0}_{n}\in\mathbb{R}^{n}$ refers to the null vector of size $n$.
  • $\ell_{1}$-norm refers to $\|\cdot\|_{1}$, $\ell_{2}$-norm refers to $\|\cdot\|_{2}$, $\ell_{1}$-loss refers to $\|\mathbf{Y}-\hat{\mathbf{Y}}\|_{1}$, and $\ell_{2}$-loss refers to $\|\mathbf{Y}-\hat{\mathbf{Y}}\|_{2}^{2}$.
@@ 1.2 Related Work
We now review the bodies of literature closest to ours. We refer interested readers to Bishop and Nasrabadi (2006) and Bishop and Bishop (2023) for a full horizon of classical ML and modern deep learning as we will not cover concepts such as backpropagation and iterative methods to optimize non-convex neural networks, e.g., stochastic gradient descent (Amari 1993).
Convex training of neural networks.
@@
Convex optimization for machine learning (ML) training is desirable as a convex training program guarantees convergence to the minimal loss. In recent years, Pilanci and Ergen (2020) have shown that training a feedforward shallow neural network (SNN), i.e., a one-hidden layer NN, with $\operatorname{ReLU}$ activation functions is equivalent to a convex training procedure. They also extended their results to convolutional neural networks (CNNs). More importantly, they fundamentally linked the $\operatorname{ReLU}$-SNN training problem to a regularized convex program. Following this work, Mishkin et al. (2022) have obtained an efficient convex training procedure of SNNs with graphics processing unit (GPU) acceleration, and Kuelbs et al. (2024) have designed an adversarial training procedure by controlling the worst-case output of the neural network in a neighbourhood around each training samples. In this work, we generalize the regularization procedure of feedforward SNNs in regression tasks through the lens of distributionally robust optimization with the Wasserstein metric. The metric is introduced in Section 2.1.
Distributionally robust neural networks.
@@
Distributionally robust optimization (DRO), in the ML setting, aims to minimize the expected training loss of an ML model over the most adverse data distribution within a distance of the training set, i.e., the distribution inside a specified ambiguity set, constructed from the available training data, that generates the biggest expected loss. As demonstrated by Kuhn et al. (2019), DRO minimizes the worst-case optimal risk associated with the training of the model. Different authors have worked on distributionally robust neural networks. Bai et al. (2023a) propose to train deep neural networks using the first-order approximation of the Wasserstein DRO loss function in the backpropagation but leave the numerical evaluation for future work. Wasserstein DRO uses the definition of the Wasserstein distance to construct the ambiguity set. We cover it in Section 2.1. The authors of Levy et al. (2020) have developed a DRO gradient estimator that integrates into neural networks’ non-convex optimization pipeline similarly to the widely used stochastic gradient descent (SGD) optimizer. In Sagawa et al. (2020), the authors propose a method to train deep NNs in a DRO fashion by defining the ambiguity set with groups of the training data and increasing regularization. These methods integrate distributional robustness in the training procedure of deep NNs but do not directly tackle the Wasserstein DRO problem, or an equivalent reformulation. They also all suffer from the lack of global optimality guarantees making them harder to deploy in large-scale critical applications. Our proposition, while only applying to shallow networks, works with the tractable reformulation of the Wasserstein DRO problem proposed by Mohajerin Esfahani and Kuhn (2018) and converges to a global optimum thanks to convexity.
Physics-informed neural networks.
The training approach for physics-informed neural networks (PINNs) primarily involves a weighted multi-objective function designed to narrow the gap between labels and model predictions while penalizing physical constraint violations. This is achieved by simultaneously minimizing the residuals of a partial differential equation (PDE) and its boundary conditions (Cuomo et al. 2022). Suppose all NN training weights are to be collected in $\bm{\Theta}$, a typical PINN training program takes the form:
$\min_{\bm{\Theta}}$ $\ \lambda_{\mathcal{F}}\mathcal{L_{F}}\left(\Phi(\mathbf{X},\bm{\Theta})\right)+\lambda_{\mathcal{B}}\mathcal{L_{B}}\left(\Phi(\mathbf{X},\bm{\Theta})\right)+\lambda_{\text{data}}\mathcal{L}\left(\Phi(\mathbf{X},\bm{\Theta}),\mathbf{Y}\right)$
@@
where $\mathcal{L_{F}}$ and $\mathcal{L_{B}}$ are the functions enforcing the differential equations and boundary conditions, respectively, of the governing physical system, $\mathcal{L}$ is the regression objective between labeled data and model predictions, and $\hat{\mathbf{Y}}=\Phi(\mathbf{X},\bm{\Theta})$ (Cuomo et al. 2022). The scalar weights $\lambda_{\mathcal{F}}$, $\lambda_{\mathcal{B}}$, and $\lambda_{\text{data}}$ must be tuned empirically. We note the lack of hard constraints. The loss terms related to the PDE are derived through the automatic differentiation of the neural network’s outputs (Baydin et al. 2017). This complex loss function is then iteratively optimized through backpropagation within the network. While authors of Cuomo et al. (2022) note that boundary conditions can be strictly imposed during training by forcing a part of the network to satisfy these conditions, see Zhu et al. (2021), they are typically integrated as penalty terms within the loss function. By using the augmented Lagrangian in the objective function and increasing the penalty terms related to constraint violations, it is possible to empirically converge to a PINN which imposes hard inequality constraints during training (Lu et al. 2021). Physics-informed neural networks are characterized by their extensive modelling capabilities, facilitated by automatic differentiation, that allows for the approximation of continuous PDEs across various neural network architectures, e.g., CNNs (Gao et al. 2021), long short-term memory networks (Zhang et al. 2020), Bayesian NNs (Yang et al. 2021), and generative adversarial networks (Yang et al. 2020). Although limited to convex constraints, in this work, we demonstrate that our training methodology can enforce hard constraints on the shallow network’s output during training without complexifying its architecture or objective function. These constraints may include, but are not limited to, boundary constraints, ramping constraints, and constraints derived from discrete-time dynamical systems.
Application.
@@
As stated previously, prediction algorithms used in large-scale VPP applications must respect some inherent constraints. They need to be computationally scalable, they must be able to model nonlinearities, and the sensibility of the application requires trustworthiness in their predictions, e.g., robustness, interpretability, and guarantees. Regarding the forecast of non-residential buildings’ energy consumption, because grid operators do not generally have access to information on building activity, occupancy, and control, the system is approached as a blackbox and data-driven approaches tend to be the default option (Amasyali and El-Gohary 2018). Authors have used support vector regression (SVR) (Dong et al. 2005) and other hybrid methods leveraging both classification and linear regression to satisfy the aforementioned constraints (Amasyali and El-Gohary 2018). These methods are linear by parts and can approximate nonlinear distributions adequately. Statistical methods such as ARIMA (Newsham and Birt 2010) and Gaussian processes have also proved to be good contenders as they even offer empirical uncertainty bounds (Weng and Rajagopal 2015). Many authors have proposed deep-learning methods, yet these propositions do not have much industrial acceptability as they fail to meet the industrial constraints and bloat the requirements for a safe ML pipeline. In this work, we show that our proposition is interesting for virtual power plant applications where the forecast of many distributed assets is used for critical decision-making as we satisfy all industrial constraints better than other comparable models.
Outlier generation.
Distributionally robust learning, in general, aims at protecting machine learning models from out-of-distribution outliers in the training set. These outliers tend to be adversarial to the prediction quality of ML models in validation and testing sets and being able to ignore them should increase models’ out-of-sample performance. As such, to correctly benchmark distributionally robust machine learning models, one must be able to generate outliers that are indeed out-of-distribution with respect to the original distribution and be able to do it regardless of the form of the original distribution. In the literature, we found that methodologies to achieve these two objectives in regression tasks are limited. For example, Chen and Paschalidis (2018) create a training set by randomly drawing samples from two distributions, the inlying and outlying one, under probability $p\in(0,1)$ and $1-p$, respectively. They used Gaussian distributions for both. In Qi et al. (2022), authors task their model with the forecast of a real-world dataset. To control the adversity of the setting, they discard the real labels and create a synthetic version of them by using the prediction of a linear regression trained on the original dataset. They then tune an additive noise over the labels to simulate different levels of outlier corruption. In Liu et al. (2022), they notably propose generating corrupted synthetic data by integrating attack mechanisms into the generation to bias their dataset, i.e., selection bias and anti-causal bias. While their methodology is interesting, they must rely on multiple Gaussian distributions to design their attacks as intended. By leveraging the sliced-Wasserstein distance (Bonnotte 2013), a tractable approximation of the Wasserstein distance, our methodology generates outliers, does not rely on any assumptions of the original distribution, making it quite flexible, and ensures that outliers do not create tangible patterns.